OpenStudy (korosh23):

Calculus Question: Topic : Limits please wait for me to write my question. Thank you

OpenStudy (korosh23):

\[\lim_{x \rightarrow \infty} (x^2 +4)^{1/2} - (x^2-1)^{1/2}\] In the above it is the question. We have to rationalize since inifinity cannot be subtracted

OpenStudy (sshayer):

\[\sqrt{x^2+4}-\sqrt{x^2-1}\times \frac{ \sqrt{x^2+4} +\sqrt{x^2-1}}{ \sqrt{x^2+4}+\sqrt{x^2-1} }\]

OpenStudy (korosh23):

The answer turned to be 5/ (x^2+4)^1/2 + (x^2-1)^1/2 = 0

OpenStudy (sshayer):

put abract also in the first portion.

OpenStudy (korosh23):

why did u square it ?

OpenStudy (korosh23):

nevermind, I got it to the power of 1/2 is square root of something. you are right

OpenStudy (sshayer):

\[\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\]

OpenStudy (korosh23):

yes

OpenStudy (korosh23):

the problem is I don't know what to do after I raionalize

OpenStudy (sshayer):

write it\[\left( \sqrt{x^2+4}-\sqrt{x^2-1} \right)\]

OpenStudy (sshayer):

put \[x=\frac{ 1 }{ y }\] \[y \rightarrow0,~as~x \rightarrow \rightarrow \infty \]

OpenStudy (korosh23):

ok, but my book gave this answer. \[\frac{ 5 }{ (x^2+4)^{1/2} +(x^2-1)^{1/2} }\]

OpenStudy (korosh23):

I don't know how to get this.

OpenStudy (korosh23):

@sshayer what should we do next?

OpenStudy (sshayer):

if you want to rationalize only then that is the correct answer. if you want to find limit then proceed as i have suggested above.

OpenStudy (korosh23):

i want to find the limit

OpenStudy (sshayer):

put x=1/y as suggested above.

OpenStudy (korosh23):

\[((1/y))^2 +4 )^1/2 - ((1/y)^2-1)^1/2\]

OpenStudy (korosh23):

ok what should i do next?

OpenStudy (sshayer):

\[\frac{ 5y }{\sqrt{1+4y^2}+\sqrt{1-y^2}}\]

OpenStudy (sshayer):

after rationalizing put x=1/y

OpenStudy (korosh23):

no

OpenStudy (sshayer):

\[now~ limit ~is~y \rightarrow 0\]

OpenStudy (korosh23):

ok can u do me a favor. Show me what you mean step by step. then I will ask u for clarification of any step I don't understand. Now I am confused.

OpenStudy (sshayer):

actually you should get after you substituting i have solved above.

OpenStudy (korosh23):

I am sorry my friend. This is not the way someone explains something to someone else who does not understand a new topic. You say substitute this and then substitute the next thing for something else. If I were you, I would write all the steps since I have the answer already, and then the person who is seeking for help would ask you to explain one step that he does not understand. I am just giving a suggestion. I hope you understand what I mean.

OpenStudy (sshayer):

when you put y=0 \[\lim_{y \rightarrow 0}\frac{ 5y }{ \sqrt{1+4y^2} +\sqrt{1-y^2}}=\frac{ 5*0 }{ \sqrt{1+0}+\sqrt{1-0} }=\frac{ 0 }{ 2 }=0\]

OpenStudy (korosh23):

I don't understand how u got 5y / square root 1 + 4y^2 + square root 1-y^2

OpenStudy (sshayer):

1. your first step is rationalize. 2.second step is put x=1/y 3. find the limit as y approaches 0

OpenStudy (sshayer):

\[\frac{ 5 }{ \sqrt{\frac{ 1 }{ y^2 }+4}+\sqrt{\frac{ 1 }{ y^2 }-1} }=\frac{ 5 }{ \sqrt{\frac{ 1+4y^2 }{ y^2 }}+\sqrt{\frac{ 1-y^2 }{ y^2 }} }\] \[=\frac{ 5 }{ \frac{ \sqrt{1+4y^2} }{ y }+\frac{ \sqrt{1-y^2} }{ y } }=\frac{ 5 }{ \frac{ \sqrt{1+4y^2}+\sqrt{1-y^2 } }{ y } }\] \[=\frac{ 5y }{ \sqrt{1+4y^2}+\sqrt{1-y^2} }\]

OpenStudy (korosh23):

perfect! :)

OpenStudy (korosh23):

just my last question is: why did my book give the answer: \[\lim_{x \rightarrow \infty} \frac{ 5 }{ (x^2+4)^1/2 + (x^2-1)^1/2 } = 0\] why is there all those variables at the bottom?qq

OpenStudy (korosh23):

I mean is there any rule or law I can know to know the answer is immediately 0.

OpenStudy (sshayer):

it is not always zero. But when the limit is infinity try to bring the variable in the denominator.

OpenStudy (yumyum247):

dayumm im studying the same thing rn...ded

OpenStudy (korosh23):

@sshayer thank you for your help. you spent a long time :)